OpenStudy (anonymous):

write the equation of the line in slope-intercept from using the given info this supposed to be a table x|0|1|2|3 y|1|5|9|13

OpenStudy (anonymous):

help please!!

myininaya (myininaya):

\[m=\frac{y_1-y_2}{x_1-x_2}=\frac{5-1}{1-0}=\frac{4}{1}=4\] \[y-y_1=m(x-x_1)\] \[y-1=4(x-0)\] \[y-1=4x\] \[y=4x+1\]

OpenStudy (anonymous):

thanks